60 cars, 21 vans and 3 other vehicles park in an underground car park during the day. What is the relative frequency of cars?
1
Expert's answer
2019-03-27T11:12:33-0400

The relative frequence of cars is equal to


"\frac{60}{60+21+3}=\frac{15}{21}=0.7143"